OpenStudy (anonymous):

Find the exact value of sin(13pi/8)

OpenStudy (anonymous):

Is one of your answer choices 0.03924146?

OpenStudy (anonymous):

or 0.08893776

OpenStudy (anonymous):

@chlomo_x3 a) (-sqrt2-sqrt2)/4 b) (sqrt2+sqrt2)/4 c) (sqrt2-sqrt2)/4 d) (-sqrt2+sqrt2)/4

OpenStudy (anonymous):

Okay this will help me gimme one sec

OpenStudy (anonymous):

okay thank you and i have two others... use the squared identities to simplify sin^2xsin^2x a) 3+4cos(2x)+cos(4x)/8 b) 3-4cos(2x)+cos(4x)/8 c) 3+4cos(2x)-cos(4x)/8 d) 3-4cos(2x)-cos(4x)/8 if cosx=2/3 and x is in quadrant 4, then sinx/2=_____ a) -1/3 b)1/3 c)-sqrt1/6 d)sqrt1/6

OpenStudy (anonymous):

sin(13pi/8) = -sin(16pi/8 - 13pi/8) = -sin(3pi/8) = -cos(4pi/8 - 3pi/8) = - cos(pi/8). Use half angle formula for cos(x): cos^2(x/2) = (cos(x) + 1) / 2 Let x = pi/4 cos^2(pi/8) = (cos(pi/4) + 1) / 2 cos^2(pi/8) = (sqrt(2) / 2 + 1) / 2 cos(pi/8) = sqrt(sqrt(2) / 4 + 1/2) -cos(pi/8) = -sqrt((sqrt(2) + 2) / 4)

OpenStudy (anonymous):

So the answer would be letter B

OpenStudy (anonymous):

The second one 2(cos^2x) = 2[(1+cos2x)/2]^2 1/2 (1+2cos2x+cos^2(2x)=1/2(1+2cos2x+1+cos4x/2 =1/4(2=4cos2x+1+cos4x)= 3+4cos2x+cos4x/4 so letter A

OpenStudy (anonymous):

thank you so much can you help with the last one @chlomo_x3